My instructor wants to make a wall chart of a computer desktop for this class.She would like the desktop icons to be proportional to the size of her 6x6 chart. if the icons are 3/8 by 3/8, how large should be on her wall chart.
Answer by Theo(13342)

you take a ratio of the area of the icon compared to the area of the computer and then apply it to the wall chart.

The wall chart is 6 feet by 6 feet.

Conver that to inches and the wall chart is 72 inches by 72 inches = 5184 square inches.

each icon on the computer is 3/8 * 3/8 = .140625 square inches.

Not sure what size computer you are using, but my computer screen measures approximately 13 inches across and 8 inches up.

That make the area of my screen 8 * 13 = 104 square inches.

The ratio of the area of an icon of .140625 square inches to the area of a typical computer screen (assuming mine is typical - yours might be different) would be .140625 / 104 = .001352163.

Apply that ratio to a screen size of 5184 square inches and you get .001352163 * 5184 = 7.009615385.

Take the square root of that to get 2.647567824 inches by 2.647567824 inches.

Round that off and your icons should be about 2.65 inches by 2.65 inches.

Let's see how that would work out.

The wall chart would hold approximately 5184 / (2.65)^2 icons which would make about 752 icons.

My computer would hold about 104 / (3/8)^2 icons which would make about 739 icons.

That's pretty close.

I would say make your icons about 2.65 inches by 2.65 inches and you would be in the ball park.

Anything between 2.5 and 3 inches by 2.5 and 3 inches should be close enough.